# Kolmogorov-Arnold Energy Models: Fast and Interpretable Generative Modeling

### v6

@@ A.4.5 Deriving the Thermodynamic Integral
$P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t)=\frac{P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)^{t}\,\,P(\bm{z}\mid\bm{f})}{\bm{Z}_{t}}$ (45)
@@
$\bm{Z}_{t}=\mathbb{E}_{P(\bm{z}\mid\bm{f})}\left[P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)^{t}\right]$ (46)
@@
$\frac{\partial}{\partial t}\log(\bm{Z}_{t})$ $=\frac{1}{\bm{Z}_{t}}\frac{\partial}{\partial t}\bm{Z}_{t}=\frac{1}{\bm{Z}_{t}}\mathbb{E}_{P(\bm{z}\mid\bm{f})}\left[\frac{\partial}{\partial t}P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)^{t}\right]$
$=\frac{1}{\bm{Z}_{t}}\mathbb{E}_{P(\bm{z}\mid\bm{f})}\left[\log P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\cdot P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)^{t}\right]\$
$=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t)}\left[\log P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right]$ (47)
@@
$\int_{0}^{1}\frac{\partial}{\partial t}\log(\bm{Z}_{t})\,dt=\int_{0}^{1}\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t)}\left[\log P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right]\,dt$ (48)
Using the fundamental theorem of calculus, (Larson & Edwards, 2008):
$\int_{0}^{1}\frac{\partial}{\partial t}\log(\bm{Z}_{t})\,dt=\log(\bm{Z}_{t=1})-\log(\bm{Z}_{t=0})=\int_{0}^{1}\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t)}\left[\log P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right]\,dt$ (49)
@@
$\log(\bm{Z}_{t=0})$ $=\log\Big(\mathbb{E}_{P(\bm{z}\mid\bm{f})}\left[P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)^{0}\right]\Big)=\log(1)=0$
$\log(\bm{Z}_{t=1})$ $=\log\Big(\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})}\left[\log P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right]\Big)=\log P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})$ (50)
@@
$\log(\bm{Z}_{t=1})-\log(\bm{Z}_{t=0})=\log P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})=\int_{0}^{1}\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t)}\left[\log P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right]\,dt$ (51)
@@ A.4.6 Discretizing the Thermodynamic Integral
Here, $t_{k}$ denotes the tempering at the $k$-th index of the schedule, and $N_{\text{t}}$ is the number of temperatures. As derived by Calderhead & Girolami (2009), Eq. 29 then be evaluated using:
$\log\left(P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})\right)=\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})+\frac{1}{2}\sum_{k=1}^{N_{t}}D_{\text{KL}}(P_{t_{k-1}}||P_{t_{k}})-D_{\text{KL}}(P_{t_{k}}||P_{t_{k-1}})$ (52)
@@
$\Delta t_{k}$ $=t_{k}-t_{k-1},\quad\quad E_{k}=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})}\left[\log\left(P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right)\right],$
$D_{\text{KL}}(P_{t_{k-1}}||P_{t_{k}})$ $=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k-1})}\Bigg[\log\Bigg(\frac{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k-1})}{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})}\Bigg)\Bigg]$ (53)
@@ A.4.7 Learning gradient for discretized Thermodynamic Integral and power posterior partition function
$D_{\text{KL}}(P_{t_{k-1}}\|P_{t_{k}})$ $=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k-1})}\Big[\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k-1})-\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})\Big]$
$\text{and}\quad D_{\text{KL}}(P_{t_{k}}\|P_{t_{k-1}})$ $=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})}\Big[\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})-\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k-1})\Big]$
$=-\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})}\Big[\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k-1})-\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})\Big]$ (54)
@@
$\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k-1})$ $-\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})=-\Delta t_{k}\cdot\log P(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi})+\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)$ (55)
@@
$\therefore D_{\text{KL}}(P_{t_{k-1}}\|P_{t_{k}})$ $=-\Delta t_{k}\cdot\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k-1})}\Big[\log P(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi})\Big]+\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)$
@@
$\text{and}\quad D_{\text{KL}}(P_{t_{k}}\|P_{t_{k-1}})$ $=\Delta t_{k}\cdot\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})}\Big[\log P(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi})\Big]-\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)$
@@
$\frac{1}{2}\sum_{k=1}^{N_{t}}D_{\text{KL}}(P_{t_{k-1}}\|P_{t_{k}})$ $-D_{\text{KL}}(P_{t_{k}}\|P_{t_{k-1}})=-\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})-2\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)$
$\therefore\log\left(P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})\right)$ $=\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})+\frac{1}{2}\sum_{k=1}^{N_{t}}D_{\text{KL}}(P_{t_{k-1}}||P_{t_{k}})-D_{\text{KL}}(P_{t_{k}}||P_{t_{k-1}})$
@@
$\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})}\Big[\nabla_{\bm{f},\bm{\Phi}}\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})\Big]$
$=\,$ $\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})}\Big[t_{k}\cdot\nabla_{\bm{\Phi}}\log P(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi})+\nabla_{\bm{f}}\log P(\bm{z}\mid\bm{f})\Big]-\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{k}})=0$
$\therefore$ $\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{k}})=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})}\Big[t_{k}\cdot\nabla_{\bm{\Phi}}\log P(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi})+\nabla_{\bm{f}}\log P(\bm{z}\mid\bm{f})\Big]$ (59)
@@
$\nabla_{\bm{f},\bm{\Phi}}\log\left(P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})\right)$ $=\sum_{k=1}^{N_{t}}\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{k}})-\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{k-1}})$
@@
$=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})}\Bigg[\nabla_{\bm{f}}\log P\left(\bm{z}\mid\bm{f}\right)+\nabla_{\bm{\Phi}}\log P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\Bigg]$
@@
A.5 Hyperparameters and training details

### v7

@@ A.4.5 Deriving the Thermodynamic Integral
$p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t)=\frac{p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)^{t}\,\,p_{\bm{f}}(\bm{z})}{\bm{Z}_{t}}$ (43)
@@
$\bm{Z}_{t}=\mathbb{E}_{p_{\bm{f}}(\bm{z})}\left[p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)^{t}\right]$ (44)
@@
$\frac{\partial}{\partial t}\log(\bm{Z}_{t})$ $=\frac{1}{\bm{Z}_{t}}\frac{\partial}{\partial t}\bm{Z}_{t}=\frac{1}{\bm{Z}_{t}}\mathbb{E}_{p_{\bm{f}}(\bm{z})}\left[\frac{\partial}{\partial t}p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)^{t}\right]$
$=\frac{1}{\bm{Z}_{t}}\mathbb{E}_{p_{\bm{f}}(\bm{z})}\left[\log p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\cdot p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)^{t}\right]\$
$=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t)}\left[\log p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right]$ (45)
@@
$\int_{0}^{1}\frac{\partial}{\partial t}\log(\bm{Z}_{t})\,dt=\int_{0}^{1}\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t)}\left[\log p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right]\,dt$ (46)
Using the fundamental theorem of calculus, (larson2008calculus):
$\int_{0}^{1}\frac{\partial}{\partial t}\log(\bm{Z}_{t})\,dt=\log(\bm{Z}_{t=1})-\log(\bm{Z}_{t=0})=\int_{0}^{1}\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t)}\left[\log p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right]\,dt$ (47)
@@
$\log(\bm{Z}_{t=0})$ $=\log\Big(\mathbb{E}_{p_{\bm{f}}(\bm{z})}\left[p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)^{0}\right]\Big)=\log(1)=0$
$\log(\bm{Z}_{t=1})$ $=\log\Big(\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})}\left[\log p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right]\Big)=\log\mathbb{P}(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})$ (48)
@@
$\log(\bm{Z}_{t=1})-\log(\bm{Z}_{t=0})=\log\mathbb{P}(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})=\int_{0}^{1}\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t)}\left[\log p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right]\,dt$ (49)
@@ A.4.6 Discretizing the Thermodynamic Integral
Here, $t_{k}$ denotes the tempering at the $k$-th index of the schedule, and $N_{\text{t}}$ is the number of temperatures. As derived by CALDERHEAD20094028, Eq. 28 then be evaluated using:
$\log\left(p_{\bm{f},\bm{\Phi}}(\bm{x}^{(b)})\right)=\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})+\frac{1}{2}\sum_{k=1}^{N_{t}}D_{\text{KL}}(\mathbb{P}_{t_{k-1}}||\mathbb{P}_{t_{k}})-D_{\text{KL}}(\mathbb{P}_{t_{k}}||\mathbb{P}_{t_{k-1}})$ (50)
@@
$\Delta t_{k}$ $=t_{k}-t_{k-1},\quad\quad E_{k}=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})}\left[\log\left(p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right)\right],$
$D_{\text{KL}}(\mathbb{P}_{t_{k-1}}||\mathbb{P}_{t_{k}})$ $=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k-1})}\Bigg[\log\Bigg(\frac{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k-1})}{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})}\Bigg)\Bigg]$ (51)
@@ A.4.7 Learning gradient for discretized Thermodynamic Integral and power posterior partition function
$D_{\text{KL}}(\mathbb{P}_{t_{k-1}}\|\mathbb{P}_{t_{k}})$ $=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k-1})}\Big[\log p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k-1})-\log p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})\Big]$
$\text{and}\quad D_{\text{KL}}(\mathbb{P}_{t_{k}}\|\mathbb{P}_{t_{k-1}})$ $=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})}\Big[\log p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})-\log p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k-1})\Big]$
$=-\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})}\Big[\log p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k-1})-\log p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})\Big]$ (52)
@@
$\log p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k-1})$ $-\log p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})=-\Delta t_{k}\cdot\log p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z})+\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)$ (53)
@@
$\therefore D_{\text{KL}}(\mathbb{P}_{t_{k-1}}\|\mathbb{P}_{t_{k}})$ $=-\Delta t_{k}\cdot\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k-1})}\Big[\log p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z})\Big]+\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)$
@@
$\text{and}\quad D_{\text{KL}}(\mathbb{P}_{t_{k}}\|\mathbb{P}_{t_{k-1}})$ $=\Delta t_{k}\cdot\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})}\Big[\log p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z})\Big]-\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)$
@@
$\frac{1}{2}\sum_{k=1}^{N_{t}}D_{\text{KL}}(\mathbb{P}_{t_{k-1}}\|\mathbb{P}_{t_{k}})$ $-D_{\text{KL}}(\mathbb{P}_{t_{k}}\|\mathbb{P}_{t_{k-1}})=-\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})-2\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)$
$\therefore\log\left(p_{\bm{f},\bm{\Phi}}(\bm{x}^{(b)})\right)$ $=\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})+\frac{1}{2}\sum_{k=1}^{N_{t}}D_{\text{KL}}(\mathbb{P}_{t_{k-1}}||\mathbb{P}_{t_{k}})-D_{\text{KL}}(\mathbb{P}_{t_{k}}||\mathbb{P}_{t_{k-1}})$
@@
$\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})}\Big[\nabla_{\bm{f},\bm{\Phi}}\log p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})\Big]$
$=\,$ $\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})}\Big[t_{k}\cdot\nabla_{\bm{\Phi}}\log p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z})+\nabla_{\bm{f}}\log p_{\bm{f}}(\bm{z})\Big]-\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{k}})=0$
$\therefore$ $\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{k}})=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})}\Big[t_{k}\cdot\nabla_{\bm{\Phi}}\log p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z})+\nabla_{\bm{f}}\log p_{\bm{f}}(\bm{z})\Big]$ (57)
@@
$\nabla_{\bm{f},\bm{\Phi}}\log\left(p_{\bm{f},\bm{\Phi}}(\bm{x}^{(b)})\right)$ $=\sum_{k=1}^{N_{t}}\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{k}})-\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{k-1}})$
@@
$=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})}\Bigg[\nabla_{\bm{f}}\log p_{\bm{f}}\left(\bm{z}\right)+\nabla_{\bm{\Phi}}\log p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\Bigg]$
@@
A.5 Sampling theory
A.5.1 Importance Sampling
Importance Sampling (IS) is a Monte Carlo that enables the estimation of expectations with respect to one probability measure using samples drawn from another. For a detailed exposition, see 910e8e7d-2e70-317c-adf6-e21cd50755a1.
The Radon-Nikodym derivative
IS involves rewriting the expectation of an arbitrary function or density, $\rho(\bm{z})$, under the target measure, $\mathbb{P}(\bm{z})$, in terms of a more tractable proposal measure, $q(\bm{z})$. Specifically, we use the Radon-Nikodym derivative $\frac{d\mathbb{P}}{d\mathbb{Q}}$ to rewrite the expectation of an arbitrary function of $\bm{z}$, denoted as $\rho(\bm{z})$:
$\mathbb{E}_{\mathbb{P}(\bm{z})}[\rho(\bm{z})]=\int_{\mathcal{\bar{Z}}}\rho(\bm{z})\,\mathbb{P}(d\bm{z})=\int_{\mathcal{\bar{Z}}}\rho(\bm{z})\cdot\frac{d\mathbb{P}}{d\mathbb{Q}}\,\mathbb{Q}(d\bm{z}).$ (59)
Here, $\frac{d\mathbb{P}}{d\mathbb{Q}}$ acts as a weight that adjusts the contribution of each sample to account for the difference between the target and proposal measures. This requires that the proposal measure $\mathbb{Q}$ is absolutely continuous with respect to the target measure $\mathbb{P}$, i.e. $\mathbb{Q}(A)=0$ implies $\mathbb{P}(A)=0$, for every measurable subset $A\in\mathcal{Z}$.
In the context of KAEM, the target distribution is the posterior, $p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})$, which is proportional to the product of the likelihood and the prior:
$p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})\propto p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z})\cdot p_{\bm{f}}(\bm{z}).$ (60)
The exact normalization constant, (the marginal likelihood $p_{\bm{f},\bm{\Phi}}(\bm{x}^{(b)})$), is intractable, but Importance Sampling circumvents the need to compute it explicitly. Instead, the expectation of $\rho(\bm{z})$ under the posterior can be expressed as:
$\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})}[\rho(\bm{z})]=\int_{\mathcal{\bar{Z}}}\rho(\bm{z})w(\bm{z})\,\mathbb{Q}(d\bm{z}),$ (61)
where the importance weights $w(\bm{z})$ are given by:
$w(\bm{z})=\frac{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})}{q(\bm{z})}=\frac{p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z})\cdot p_{\bm{f}}(\bm{z})}{q(\bm{z})}.$ (62)
Proposal distribution and importance weights
A practical choice for the proposal density $q(\bm{z})$ is the prior, $p_{\bm{f}}(\bm{z})$, given the availability of draws from the prior using the method outlined in Sec. 2.4.3. This simplifies the importance weights, as the prior has a tractable form and covers the support of the posterior. This yields:
$w(\bm{z}^{(s)})=p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z}^{(s)})\propto\exp(\log(p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z}^{(s)}))).$ (63)
$\frac{w(\bm{z}^{(s)})}{\sum_{r=1}^{N_{s}}w(\bm{z}^{(r)})}=\text{softmax}_{s}\Big(\log(p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z}^{(s)}))\Big).$ (64)
Resampling
To address this, the weights are resampled before proceeding with the posterior expectation. This is accomplished by using one of the methods outlined by douc2005comparisonresamplingschemesparticle, whenever the ESS falls below a certain threshold determined by Eq. 26. Resampling redistributes the sample weights, creating a new population with uniformly distributed weights while preserving the statistical properties of the original distribution. As training progresses and the prior becomes a better match for the posterior, the need for resampling decreases.
After resampling, the corresponding weights will be uniformly distributed and a standard Monte Carlo estimator can be applied instead of the IS estimator.
Importance Sampling estimator
$\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})}[\rho(\bm{z})]$ $\approx\sum_{s=1}^{N_{s}}\rho(\bm{z}^{(s)})\cdot w_{\text{norm}}(\bm{\bm{z}^{(s)})}$ (65)
$\approx\frac{1}{N_{s}}\sum_{s=1}^{N_{s}}\rho(\bm{z}^{(s)}_{\text{resampled}})$. (66)
The estimator is unbiased, implying that its expectation matches the true posterior expectation. As $N_{s}\to\infty$, the approximation converges almost surely to the true value, provided that $q(\bm{z})$ sufficiently covers the posterior to ensure a well-conditioned Radon-Nikodym derivative. This estimator is used for the log-marginal likelihood and its gradient in Eqs. 15 &  16, since it provides a means of estimating expectations with respect to the posterior distribution.
A.5.2 Residual resampling
In this study, we adopted residual resampling to filter the population of latent samples whenever the effective sample size (ESS) criterion in Eq. 26 is satisfied. Given a set of normalized weights $\{w_{\text{norm}}(\bm{z}^{(s)})\}_{s=1}^{N}$, the method deterministically replicates samples according to the integer part of $Nw_{\text{norm}}$, and uses multinomial sampling to allocate the remaining draws according to the residual mass.
  1. 1.
    Integer replication: Each sample $s$ is deterministically replicated
    $r_{s}=\left\lfloor N\,w_{\text{norm}}(\bm{z}^{(s)})\right\rfloor,$ (67)
    resulting in $\sum_{s=1}^{N}r_{s}$ allocated indices.
  2. 2.
    Residual weights: The number of remaining draws is
    $R=N-\sum_{s=1}^{N}r_{s}.$ (68)
    The unallocated probability mass (fractional remainder) is computed as
    $m_{s}=N\,w_{\text{norm}}(\bm{z}^{(s)})-r_{s}.$ (69)
    If $R>0$, the residual weights are obtained by normalizing this mass:
    $w_{\text{residual}}(\bm{z}^{(s)})=\frac{m_{s}}{R},$ (70)
    otherwise $w_{\text{residual}}(\bm{z}^{(s)})=0$ for all $s$.
  3. 3.
    Multinomial allocation: The remaining $R$ indices are sampled according to the cumulative distribution function of the residual weights. For each draw,
    $k^{*}=\min\left\{j\mid\sum^{j}_{s=1}w_{\text{residual}}(\bm{z}^{(s)})\geq u_{i}\right\}\quad\text{where}\quad u_{i}\sim\mathcal{U}([0,1]).$ (71)
    The final resampled index set consists of the deterministic replications given by $\{r_{s}\}$ and the stochastic residual allocations drawn from Eq. 71. After resampling, weights are reset to a uniform distribution and a standard Monte Carlo estimator can be used.
A.6 Hyperparameters and training details